# Implicit Actor Critic Coupling via a Supervised Learning Framework for RLVR

### v1

@@ 3.4 Score Function Instantiation
To address the potential distributional imbalance in generated outputs $o_{i}$ for $q$, we adopt the class imbalance treatment methodology proposed by (King & Zeng, 2001), whereby differential weights are assigned to correct and incorrect samples respectively. Specifically, when the proportion of correct and incorrect samples in the training data exhibits imbalance, we mitigate the adverse effects of sample distribution bias on model performance by adjusting weight parameters to balance the model’s attention across different categorical samples.
MATH 500 ($\text{pass@}k$) AMC 23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 56.66 67.64 75.75 81.91 86.87 90.60 30.20 66.16 83.88 90.42 94.85 97.10
PPO 65.48 71.94 77.51 82.14 86.02 89.20 38.57 66.07 80.19 86.81 92.09 95.53
GRPO 66.92 73.34 78.60 83.01 86.60 89.00 37.01 68.71 82.61 87.72 92.39 96.44
PACS 67.31 74.75 80.60 85.03 88.22 90.40 39.63 68.23 82.45 88.23 92.70 96.18
AIME 2024 ($\text{pass@}k$) AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 4.92 15.66 27.21 33.45 40.95 49.68 1.73 10.62 23.45 30.69 39.09 47.38
PPO 7.55 19.70 30.27 35.69 42.08 50.81 1.40 8.99 21.11 28.07 35.52 42.51
GRPO 8.14 19.13 26.06 31.65 37.75 42.76 2.62 7.07 14.07 20.18 28.37 36.58
PACS 7.98 21.49 30.10 35.87 42.25 48.32 2.94 11.32 22.91 30.50 39.07 47.18
Table 1: Results of Qwen2.5-3B trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best. Due to space constraints, selected results are presented here. The complete results are provided in AMC23 (Table 5), AIME 2024 (Table 7) and AIME 2025 (Table 9) in the Appendix.
@@ 4.1 Experimental Setup
This study uniformly adopts the DeepScaleR (Luo et al., 2025) dataset as the training corpus, which constitutes a high-quality mathematical problem-solving collection comprising approximately 40,000 unique mathematical problem-answer pairs. To comprehensively evaluate the performance of methods, we conduct extensive evaluations on four representative benchmarks: MATH 500, AMC23, AIME 2024, and AIME 2025. The experiments utilize Qwen2.5-3B and Qwen2.5-7B (Qwen et al., 2025) as base models, enabling us to assess PACS’s performance across various model scales.
Baselines & Hyperparameters.
@@
To evaluate the effectiveness of our approach, we compare the performance of PACS with representative RL algorithms: PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). The training objectives of these algorithms can be found in Appendix. We also report the performance of base models for reference. The train batch size is set to 1,024. The learning rate of actor model is set to 1e-6. For each query, $8$ responses are sampled. For PACS, $\beta$ is set to 1.0. During inference, we configure the sampling parameters with a temperature of 0.6 and top-$p$ of 0.96. For Qwen2.5-3B, we conduct training for 140 steps, while for Qwen2.5-7B, we extend the training to 300 steps. Prompt template and detailed hyperparameter configuration can be found in the Appendix.
Hardware.
@@
All experiments are conducted on NVIDIA H200 GPUs. Training employs verl (Sheng et al., 2024), an efficient reinforcement learning library specifically designed for large language models. For inference, we utilize vllm (Kwon et al., 2023).
Evaluation Metrics.
To evaluate the model’s capability to generate diverse correct solutions while mitigating the randomness bias that may arise from single sampling, we employ the $\text{pass@}k$ metric for assessment. $\text{pass@}k$ represents the probability that at least one correct answer exists among $k$ generated candidate solutions. $\text{pass@}k$ also enables performance evaluation under different computational budgets by adjusting the value of $k$. However, directly calculating $\text{pass@}k$ with only $k$ samples per instance often results in high variance. To obtain more stable evaluation results, we generate $n$ samples for each problem (where $n\geq k$), count the number of correct samples $c$, and calculate the unbiased estimator (Chen et al., 2021):
$\text{pass@}k=\mathbb{E}_{x\sim D}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]. (10)
Our sampling strategy adapts to the scale of dataset. For the MATH 500 dataset, given its scale of 500 problems, the sampling size is set to $n=32$ for each problem and model performance is evaluated across $k\in\{1,2,4,8,16,32\}$. Conversely, for smaller test sets (AMC23, AIME 2024, and AIME 2025), which contain fewer problems but with high-quality content, larger sampling scale is employed to obtain more stable and reliable statistical results. Specifically, we generate 512 candidate solutions for each problem ($n=512$) and conduct evaluation across a broader range of $k$ where $k\in\{1,2,4,8,16,32,64,128,256\}$.
MATH 500 ($\text{pass@}k$) AMC 23 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 61.55 73.90 82.09 87.48 90.93 92.80 35.72 70.21 85.53 91.19 95.89 99.18
PPO 76.56 80.74 84.15 86.70 88.60 90.20 51.22 72.80 81.92 85.71 88.67 90.92
GRPO 78.29 82.08 84.70 86.60 88.26 90.00 57.22 72.20 80.72 86.14 90.69 93.34
PACS 78.97 84.39 88.18 90.68 92.33 93.60 57.83 78.17 88.22 91.95 95.51 98.51
    - w/o weight 77.84 83.74 87.68 90.30 92.06 93.20 54.51 75.15 83.77 86.89 88.97 90.62
AIME 2024 ($\text{pass@}k$) AIME 2025 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$ $\bm{k=1}$ $\bm{8}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
Base 7.26 23.46 36.41 43.77 52.66 62.55 3.35 16.04 28.56 34.99 42.07 50.11
PPO 15.56 28.50 38.71 41.50 43.85 46.46 9.61 24.89 32.32 34.94 38.27 42.91
GRPO 14.10 25.37 34.78 38.50 41.66 45.42 8.36 21.63 33.26 38.12 42.33 46.45
PACS 15.80 29.75 43.90 51.21 56.54 59.78 13.33 29.65 42.88 49.45 54.48 58.22
    - w/o weight 15.72 24.57 32.58 38.77 46.16 53.42 10.44 22.86 31.03 36.59 42.39 47.67
@@
Table 2: Results of Qwen2.5-7B trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best. Due to space constraints, selected results are presented here. The complete results are provided in AMC23 (Table 6), AIME 2024 (Table 8) and AIME 2025 (Table 10) in the Appendix.
Figure 3: Performance analysis of PACS with varying $\beta$. The 3D heatmaps show pass@k scores for different combinations of $\beta$ values (0.1, 0.5, 1, 2, 10) and $k$ values on MATH-500, AMC23, AIME-2024, and AIME-2025 datasets.
@@ 4.2 Main Results
Table 1 and Table 2 compare the overall results of PACS with different baselines. Experimental results on the Qwen2.5-3B model demonstrate that PACS exhibits significant performance advantages across multiple mathematical reasoning tasks. On the MATH 500 dataset, PACS achieves optimal performance across all $k$-value configurations, particularly attaining a pass rate of 67.31% at $\text{pass@}1$, representing an improvement of 10.65 points over the base model and surpassing both PPO and GRPO. In the AMC 23 task, PACS similarly demonstrates exceptional performance. Notably, on the more challenging AIME 2024 and AIME 2025, PACS maintains its leading position across most $k$-value settings, demonstrating the algorithm’s stability and effectiveness in complex mathematical reasoning tasks.
On the larger-scale Qwen2.5-7B model, the advantages of the PACS algorithm become even more pronounced, further validating its generalization capability across different model scales. On the MATH 500 dataset, PACS achieves superior performance across all evaluation metrics. The performance of PACS is especially remarkable on the AIME series tasks: in AIME 2024, it achieves a pass rate of 59.78% at $\text{pass@}256$, representing improvements of 13.32 and 14.36 points compared to PPO and GRPO, respectively; in AIME 2025, the $\text{pass@}256$ metric shows improvements of 15.31 and 11.77 points compared to PPO and GRPO, respectively. These results conclusively demonstrate PACS’s exceptional performance in handling high-difficulty mathematical reasoning problems and its significant advantages over reinforcement learning methods PPO and GRPO.
4.3 Ablation Study
Different $\beta$.
To validate the impact of $\beta$ on the model’s performance in Equation 8, we design and conduct ablation experiments. Models are trained with different $\beta\in\{0.1,0.5,1,2,10\}$ and their performance is evaluated in Figure 3. The results demonstrate that model achieves optimal performance when $\beta=1$, yielding higher $\text{pass@}k$.
Further analysis reveals that the AIME 2024 and AIME 2025 datasets exhibit heightened sensitivity to variations in $\beta$. This suggests that for tasks of higher complexity and difficulty, $\beta$ constitutes a critical step in achieving optimal performance. In contrast, the AMC23 and MATH 500 datasets demonstrate greater robustness across different $\beta$, with minor fluctuations in model performance. This phenomenon can be attributed to the relatively simpler nature of problems in these datasets, enabling the model to maintain stable performance across a broader range of hyperparameters.
Different Advantage Estimators.
To investigate the influence of advantage estimators on PACS’s efficacy, we extend our evaluation beyond the default RLOO to include GRPO (Shao et al., 2024) and Dr. GRPO (Liu et al., 2025b) as shown in Appendix C.4. Our findings indicate that the PACS with RLOO exhibits superior overall performance, a lead that becomes more significant at higher $\text{pass@}k$ metrics. Notably, whereas RLOO and Dr. GRPO yield comparable results on datasets such as MATH 500 and AMC23, RLOO’s superiority is markedly evident on the more demanding AIME benchmark. We posit that this performance gap stems from a subtle algorithmic divergence between RLOO and Dr. GRPO, which is magnified under more challenging conditions. By contrasting a single sample with the cohort average, RLOO’s leave-one-out mechanism yields more robust and precise signals for credit assignment.
Effect of Weight Integration.
To validate the effectiveness of the weight mechanism in PACS, we conduct ablation experiments by comparing the performance of PACS with and without the weight component. The results presented in Table 2 clearly demonstrate the positive impact of the weight mechanism. PACS consistently outperforms its weight-free variant (w/o weight). The benefits are particularly pronounced on challenging AIME datasets, where the weight component proves crucial for complex mathematical reasoning tasks, achieving substantial performance improvements especially at higher sampling rates (e.g., 6.36% point improvement on AIME 2024 at $\text{pass@}32$ of Qwen2.5-7B).
4.4 Training Dynamics
Figure 4: Training dynamics of Qwen2.5-7B on the DeepScaleR dataset across different optimization methods (PPO, GRPO, PACS, and PACS w/o weight). The plots illustrate the evolution of three key metrics during training: (a) entropy loss, (b) gradient norm, and (c) mean response length over training steps.
Figure 4 compares training dynamics of four methods applied to Qwen2.5-7B. PPO and GRPO maintain consistently low entropy loss, resulting in entropy collapse and conservative outputs. But PACS shows a non-monotonic pattern: entropy increases to about 1,700 at step 220, then stabilizes at 1,200-1,300, reflecting a superior exploration-exploitation balance. PACS w/o weight demonstrates intermediate behavior. Gradient norms reveal similar patterns. While PPO and GRPO maintain stable gradient norms throughout training, PACS peaks at 40 with sustained elevation, enabling aggressive parameter updates and broader exploration. This sustained gradient elevation in PACS, contrasting with the stable but lower gradient norms in baseline methods, demonstrates superior training efficiency. The consistently higher gradient activity indicates that PACS continues making meaningful parameter updates throughout the training process, while the stable but modest gradients in PPO and GRPO suggest more conservative optimization that may lead to premature convergence. Response lengths show PACS generating the most detailed responses. PACS achieves superior exploration capability and response quality through moderate entropy maintenance and dynamic gradient updates, providing novel optimization insights compared to other methods.

### v2

@@ 3.5 Practical Considerations: Handling Data Imbalance
To address the potential distributional imbalance in generated outputs $o_{i}$ for $q$, we adopt the class imbalance treatment methodology proposed by (King and Zeng, 2001), whereby differential weights are assigned to correct and incorrect samples respectively. Specifically, when the proportion of correct and incorrect samples in the training data exhibits imbalance, we mitigate the adverse effects of sample distribution bias on model performance by adjusting weight parameters to balance the model’s attention across different categorical samples.
@@ 4.1 Experimental Setup
Datasets & Models. We utilize the DeepScaleR dataset as the training corpus (Luo et al., 2025), which constitutes a high-quality mathematical problem-solving collection. For evaluation, we conduct extensive evaluations on five representative benchmarks: MATH 500, AMC 23, AIME 2024, AIME 2025 and BeyondAIME. The experiments utilize Qwen3-4B and Qwen3-8B (Qwen et al., 2025) as base models, assessing PACS across different model scales.
@@
Baselines & Hyperparameters. To comprehensively evaluate the effectiveness of PACS, we compare PACS with a diverse set of baselines categorized into three groups: 1). Closed-source models: including state-of-the art models such as Claude Opus 4.1 and Gemini 2.5 Pro (Comanici et al., 2025), whose results are directly cited from prior works (Team et al., 2025a; Li et al., 2025); 2). Open-source models: including recent models utilizing advanced post-training of RL techniques, such as Qwen3 models (Yang et al., 2025), Deepthought-8B (Ruliad, 2024), Eurus-2-7B-PRIME (Cui et al., 2025), LIMO (Ye et al., 2025), Marco-o1 (Zhao et al., 2024), OpenThinker3-7B (Guha et al., 2025) and Sky-T1-7B (Team, 2025a); and 3). Representative RL algorithms: we implement and train PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) on the same base models to ensure a fair comparison. The training objectives of these algorithms can be found in Appendix B. We also report the performance of base models for reference. The train batch size is set to 1,024. For each query during training, $8$ responses are sampled. For PACS, $\beta$ is set to 1.0. During inference, we configure the sampling parameters with a temperature of 0.6. Prompt template and detailed hyperparameter configuration can be found in the Appendix C.
@@
Hardware. All experiments are conducted on NVIDIA H200 GPUs. We employ verl (Sheng et al., 2024), an reinforcement learning library specifically designed for LLMs. For inference, we utilize vllm (Kwon et al., 2023).
[…]
AMC 23 ($\text{pass@}k$) AIME 2024 ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
Base 81.74 90.53 92.31 94.00 95.65 96.86 46.98 59.78 65.10 69.28 72.12 74.99
PPO 86.27 92.36 93.91 95.46 96.70 98.13 54.61 71.46 75.86 79.11 81.75 83.12
GRPO 87.83 93.17 94.71 95.92 97.91 99.69 54.56 74.52 80.08 82.42 83.21 83.33
PACS 88.69 94.28 95.36 96.03 96.72 97.35 57.58 76.24 81.21 83.30 84.14 85.00
   - w/o weight 89.63 94.80 95.53 96.02 96.50 97.32 58.93 76.64 81.13 83.21 83.59 84.73
AIME 2025 ($\text{pass@}k$) BeyondAIME ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
Base 33.83 45.86 51.19 56.12 60.54 65.74 16.28 23.00 26.75 31.36 36.65 41.74
PPO 42.42 55.67 61.70 68.25 73.99 77.43 23.38 33.60 38.83 44.10 49.11 53.80
GRPO 45.29 58.80 64.04 69.12 73.84 77.39 25.79 37.61 43.08 48.40 53.32 57.64
PACS 46.38 61.93 67.78 72.20 75.71 79.12 28.86 43.15 49.58 55.39 60.19 64.02
   - w/o weight 46.15 60.07 66.29 71.98 76.10 78.23 28.52 40.43 45.55 50.10 54.35 58.97
@@
Table 2: Results of Qwen3-8B trained with PPO, GRPO and PACS. Bold numbers indicate the best performance. Underlined numbers indicate the second best. Due to space constraints, results on MATH 500 are shown in Table 8.
@@ 4.2 Main Results
Comparative Analysis with Existing Models. table 1 shows that PACS significantly enhances the reasoning capabilities of base models across all benchmarks. Specifically, PACS-8B achieves an average $\text{pass@}1$ of 63.32%, surpassing Qwen3-8B by 9.57 points and outperforming strong baselines like LIMO (52.05%). Notably, PACS exhibits exceptional parameter efficiency: PACS-4B (62.63%) not only beats 7B-scale competitors but also outperforms the significantly larger Qwen3-14B on challenging tasks like AIME 2025 (45.63% vs. 36.22%). The gains are particularly pronounced on complex benchmarks; for instance, on AIME 2025 and BeyondAIME, PACS-8B improves upon the base model by over 12 points, effectively unlocking deep chain-of-thought capabilities.
Superior Performance against RL Baselines. Comparing algorithmic efficacy in tables 2 and 6, PACS consistently outperforms PPO and establishes a decisive advantage over GRPO on complex reasoning tasks. This superiority is most evident on the hardest benchmark, BeyondAIME. Here, PACS-8B achieves a $\text{pass@}64$ of 64.02%, surpassing GRPO (57.64%) by over 6 points, while PACS-4B similarly improves the metric from 57.58% to 62.10%. Furthermore, PACS-8B maintains a comprehensive lead over GRPO on AIME 2024 and 2025 in both single- and multi-sample settings. These results indicate that PACS’s supervised learning formulation, which implicitly couples actor and critic updates, provides more effective optimization signals than pure RL approaches for navigating complex solution spaces.
Training Dynamics and Performance Analysis. To further investigate the optimization efficiency, we visualize the evaluation trajectories of Qwen3-8B throughout the training process in Figure 3. The curves report $\text{pass@}1$ on benchmarks, evaluated every 20 training steps. The results highlight PACS’s robustness across diverse problem distributions. On benchmarks like AMC 23 and the highly challenging BeyondAIME, PACS exhibits high training efficiency, rapidly establishing a decisive performance lead over RL baselines. Notably, on AIME 2025, the performance gap narrows as GRPO also demonstrates competitive learning trajectories; yet, PACS remains resilient, matching the strong baseline’s pace and securing the best final performance.
Figure 3: Training dynamics of Qwen3-8B. The curves illustrate the evolution of $\text{pass}@1$ on math benchmarks throughout the training process.
5 Ablation Study
5.1 Different $\beta$
Figure 4: Performance analysis of PACS with varying $\beta$. The 3D heatmaps show $\text{pass@}k$ scores for different combinations of $\beta$ values (0.1, 0.5, 1, 2, 10) and $k$ values on AMC23, AIME-2024, AIME-2025 and BeyondAIME.
To validate the impact of $\beta$ on the model’s performance in eq. 7, we design and conduct ablation experiments. Models are trained with different $\beta\in\{0.1,0.5,1,2,10\}$ and their performance is evaluated in fig. 4. The results demonstrate that PACS is highly robust to variations in $\beta$. While extreme values (e.g., $\beta=10$) maintain competitive performance on easier tasks like AMC 23, a distinct ”sweet pot” emerges in $\beta\in[0.5,1.0]$ for maximizing $\text{pass@}1$ on complex reasoning tasks. Given that $\text{pass@}1$ is often the primary objective for reasoning models, we select $\beta=1$ as the optimal setting, effectively balancing the reward signal strength to achieve superior reasoning performance.
5.2 Different Score Function
To validate the effectiveness of our scoring mechanism, we compare the RLOO score function against GRPO (Shao et al., 2024) and Dr. GRPO (Liu et al., 2025c). Detailed results are shown in Appendix D.3. Our analysis reveals that PACS with RLOO generally yields superior performance, particularly on complex reasoning tasks. For instance, on BeyondAIME benchmark, PACS achieves a $\text{pass@}1$ of 27.16%, consistently outperforming both GRPO (26.41%) and Dr. GRPO (26.67%). We attribute this robustness to the leave-one-out mechanism of RLOO, which provides unbiased advantage estimates with lower variance compared to the group-mean baselines used in GRPO variants.
5.3 Effect of Weight Integration
To validate the effectiveness of the weight mechanism in PACS, we conduct ablation experiments by comparing the performance of PACS with and without the weight component as shown in tables 2, 6, 7 and 8. For Qwen3-4B, PACS systematically outperforms the unweighted variant across benchmarks; notably, it achieves a pass@1 of 90.45% on AMC 23 and 55.10% on AIME 2024, surpassing the w/o weight PACS scores of 89.18% and 54.56% respectively. This advantage is further amplified on complex tasks with the larger Qwen3-8B model. While performance on simpler tasks remains comparable, PACS establishes a clear lead on the hardest benchmark, BeyondAIME. It not only improves the $\text{pass@}1$ accuracy to 28.86% compared to 28.52% , but more importantly, achieves a remarkable 64.02% at pass@64, exceeding the unweighted variant’s 58.97%.
Figure 5: Exploration and Diversity Analysis. (a) Entropy loss dynamics for Qwen3-4B (top) and 8B (bottom). Unlike baselines that suffer from entropy collapse, PACS maintains higher entropy, enabling sustained exploration.(b) Centered PCA projection of correct solutions for sampled problems (ID shown in bottom-right). The broader semantic coverage of PACS (Blue) compared to GRPO (Red) visually confirms superior diversity.
6 Generalization and Diversity Analysis
In this section, we analyze the broader capabilities of PACS by first evaluating its generalization to out-of-domain tasks to rule out overfitting, and then exploring the solution diversity that underpin this mechanism.
6.1 Generalization of Out-of-Domain Tasks
Figure 6: Performance on out-of-domain tasks. The chart reports the evaluation scores of Qwen3-8B on LiveCodeBench and GPQA. PACS achieves substantial gains over PPO and GRPO.
A key challenge in RLVR is preventing the policy from over-specializing to the training domain, which can lead to a degradation of general capabilities. To evaluate whether PACS learns fundamental reasoning skills that transfer across domains, we assess the Qwen3-8B model trained on the DeepScaleR dataset (math) on two out-of-domain benchmarks: LiveCodeBench (Jain et al., 2024) [^2] Evaluations are performed on the LiveCodeBench dynamic subset spanning the period from 7/1/2024, through 1/1/2025. and GPQA (Rein et al., 2023). For these assessments, we utilize the toolkit provided by GLM-4.5, adhering to the default settings for all experimental configurations. As shown in Figure 6, PACS demonstrates superior cross-domain generalization compared to PPO and GRPO. On LiveCodeBench, PACS achieves 45.96, significantly outperforming the Vanilla model (31.06) as well as PPO (35.87) and GRPO (38.01), which suggests that the reasoning chains optimized by PACS effectively transfer to code logic generation while baseline RL methods yield diminishing returns. This advantage is further corroborated on GPQA, where PACS attains 52.46, surpassing both GRPO (48.55) and PPO (48.48). These results indicate that PACS does not merely memorize domain-specific templates; instead, by treating outcome rewards as supervised signals within a stable optimization framework, it enhances the model’s intrinsic reasoning capabilities.
6.2 Exploration and Diversity
Model Method AIME2024 AIME2025
Qwen3-4B PACS 0.0278 0.0300
GRPO 0.0260 0.0266
Qwen3-8B PACS 0.0272 0.0392
GRPO 0.0217 0.0249
Table 3: Comparison of diversity scores on AIME 2024 and AIME 2025. PACS consistently exhibits higher diversity than GRPO across different model scales.
We observe that PACS effectively mitigates the entropy collapse issue commonly found in RL baselines, leading to semantically richer and more diverse solutions as shown in fig. 5(a). To quantify the diversity of the generated reasoning paths, we utilize the average pairwise consine distance between the embeddings of responses. A higher diversity score ($S_{div}$) indicates greater semantic variation. The definition of $S_{div}$ is detailed in Appendix E. We assess the response diversity on two representative benchmarks: AIME 2024 and AIME 2025. We specifically compare PACS with GRPO using the Qwen3-4B and Qwen3-8B. Note that we compute $S_{div}$ on correct responses, as our goal is to achieve diverse yet valid solutions. As shown in table 3, PACS consistently surpasses GRPO in diversity (0.0392 vs. 0.0249). While the rigid nature of mathematical solutions limits absolute scores, PACS achieves a substantial relative improvement, effectively mitigating mode collapse and fostering semantically distinct reasoning paths.
To intuitively compare the diversity of reasoning paths, centered PCA approach is employed. We randomly sample some problems. fig. 5(b) displays the 2D PCA projection of the embeddings for correct solutions generated by PACS (Blue) and GRPO (Red). The data is mean-centered to align the centroids of both distributions at the origin $(0,0)$. The shaded regions (KDE contours) and scatter points illustrate that PACS covers a significantly larger area in the semantic space, indicating a more diverse set of reasoning paths compared to the more concentrated distribution of GRPO.